OpenStudy (cleblanc17):

Medaling who ever responds first solve for x 10=3/2x+1

OpenStudy (anonymous):

x=6

OpenStudy (thomas5267):

\[ 10=\frac{3}{2}x+1\\ 9=\frac{3}{2}x \] What can you do next to get rid of the 3/2?

OpenStudy (cleblanc17):

you can multiply right<>>?

OpenStudy (anonymous):

what do you mean

OpenStudy (thomas5267):

Yes you can.

OpenStudy (anonymous):

have any more

OpenStudy (cleblanc17):

no thats it... so i take the 3 and put it on the outside

OpenStudy (anonymous):

yes your trying to get the x alone so you divide 9 and 3/2

OpenStudy (cleblanc17):

ohhh

OpenStudy (cleblanc17):

then its 1.5 ....

OpenStudy (thomas5267):

I don't think x=1.5. Try put it back into the equation and you will see why.

OpenStudy (cleblanc17):

oh lol i did it wrond its 3/2 divided by 9

OpenStudy (cleblanc17):

im feeling dumber than i started out tbh ;(

OpenStudy (anonymous):

this is how you put in 9 divide 3/2

OpenStudy (anonymous):

you are smart you just got to bring it out

OpenStudy (anonymous):